OpenStudy (anonymous):

solve by substitution 2.5x-3y=-3 3.25x-y=-14

OpenStudy (anonymous):

rearrange the first equation in terms of y 2.5x - 3y=-3 2.5x + 3 = 3y y= 2.5/3x + 3/3 y=2.5/3x + 1 now substitute that as y into the second equation 3.25x-y=-14 try it yourself.

OpenStudy (anonymous):

\[2.5x-3y=-3\] Use the second equation to make y the subject. Since you just need to move values to the LHS and RHS. You don't need to do much about it. \[y=3.25x+14\] Now for every y you see in the first equation write (3.25x+14). \[2.5x-3(3.25x+14)=-3\]

OpenStudy (anonymous):

\[2.5x-3(3.25x)-3(14)=-3\] \[2.5x-10.75x-42=-3\]

OpenStudy (anonymous):

Collect like terms and solve for x.

OpenStudy (anonymous):

hold up im just working it out!

OpenStudy (anonymous):

wait wouldnt it be 9.75

OpenStudy (anonymous):

instead of 10.75 wouldnt be 9.75x

OpenStudy (anonymous):

Oh yeah. Good spotting mate. Excellent work. You should always look at everything I wrote or you might of copied wrong working out. ;)

OpenStudy (anonymous):

That's good work ethic. Always do that and you will succeed in math.

OpenStudy (anonymous):

So. ][2.5x-9.75x-42=-3\]

OpenStudy (anonymous):

\[2.5x-9.75x-42=-3\]

OpenStudy (anonymous):

haha thank you! would the answer be x=-.1858

OpenStudy (anonymous):

Nope, incorrect. Start by collecting like terms.

OpenStudy (anonymous):

Show me when you collected your like terms.

OpenStudy (anonymous):

okay so i collected 2.5x-9.75 and i got -7.25 correct?

OpenStudy (anonymous):

okay so i collected 2.5x-9.75xand i got -7.25x correct?

OpenStudy (anonymous):

Okay. I'm a bit of a hater towards negative signs on x's. So I would move the -7.25x to the RHS so that you eliminate that minus sign.

OpenStudy (anonymous):

Move the constants to the LHS.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

RHS=Right Hand Side

OpenStudy (anonymous):

\[-7.25x-42=-3\]

OpenStudy (anonymous):

Move the term \[-7.25x\] to the RHS (Right Hand Side of the equal sign.)

OpenStudy (anonymous):

So \[-42=-3+7.25x\]

OpenStudy (anonymous):

Move the -3 to the LHS(Left Hand Side of the equal sign). Could you do that for me?

OpenStudy (anonymous):

yes so the answer is -5.37

OpenStudy (anonymous):

Correct buy you didn't round up properly.

OpenStudy (anonymous):

but*

OpenStudy (anonymous):

so -5.4

OpenStudy (anonymous):

\[-5.379\neq -5.37 \]

OpenStudy (anonymous):

I mean you didn't round up to the nearest hundredth properly.

OpenStudy (anonymous):

\[-5.379=-5.38\]

OpenStudy (anonymous):

5.4 is also correct. But it depends on whether your question asks you to round up to the nearest decimal place or the nearest 2 decimal places.

OpenStudy (anonymous):

it doesnt say that so would it just be 5.4

OpenStudy (anonymous):

If your questions does not ask you to round up. Answer it as a fraction.

OpenStudy (anonymous):

it doesnt say round or anything

OpenStudy (anonymous):

question*

OpenStudy (anonymous):

and i still have to solve for y right?

OpenStudy (anonymous):

Do not round up if they don't ask you to. Always put it to the best possible answer which is a fraction. Fractions is much more accurate than the rounded up decimal.

OpenStudy (anonymous):

Yes. That's why never round up when you need to use that number to find another number.

OpenStudy (anonymous):

So substitute the fraction into one of the two equations given to you from your question (preferrably the second equation) to find the y-value.

OpenStudy (anonymous):

5.4= 27/5 as a fraction? so i would sub that in for x in the seconf equation

OpenStudy (anonymous):

No. 27/5 is still 5.4. That's still the rounded up fraction.

OpenStudy (anonymous):

And it's meant to be a negative fraction not a positive.

OpenStudy (anonymous):

\[-\frac{156}{29}\]